OpenStudy (anonymous):

how do you work out the integral of 4t^3 times squareroot(1+9t^4)dt

OpenStudy (anonymous):

try a substitution of a trig function

OpenStudy (anonymous):

a trig function? l dont follow

OpenStudy (muhammad_nauman_umair):

As, \[\int\limits4t^3*\sqrt{1+9t^4} dt=1/9(\int\limits36t^3*\sqrt{1+9t^4} dt)=1/9*{(1+9t^4)^{3/2}}/(3/2)\]

OpenStudy (muhammad_nauman_umair):

then \[2/27*(1+9t^4)^{3/2}+c \]

OpenStudy (callisto):

Sorry , a little bit messy :(

OpenStudy (muhammad_nauman_umair):

Got my way or not

OpenStudy (callisto):

@Muhammad Nauman Umair some of your steps are missing because of the usage of equation...

OpenStudy (muhammad_nauman_umair):

the steps which i skipped are not difficult to understand @callisto

OpenStudy (anonymous):

oh thanx guys that was really helpful

OpenStudy (muhammad_nauman_umair):

welcome